Find the unknown qanity for an account that earns simple annual interest.



A = $6240, P = $4000 r =____, t = 7 years
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

6240=4000(1+R)^7
6240/4000=(1+R)^7
1.56=(1+R)^7
(1+R)=7th ROOT 1.56
1+R=1.0659
R=1.0659-1.
R=.0659 OR 6.59% IS THE INTEREST RATE.